OpenStudy (anonymous):

A cone has a base area of 50 in^2 and a lateral area which is 10 times bigger than the base area. The surface area of the cone is _____ in^2 Answers: 500 550 600 650

OpenStudy (wolf1728):

Seems you don't need any formulas for this. base = 50 sq inches lateral area = base * 10 = 500 Total cone area = 50 + 500 = 550 sq inches

OpenStudy (anonymous):

Thank you!! :)

OpenStudy (wolf1728):

u r welcome :-)